# A Systematic Survey on Large Language Models for Algorithm Design

### v2

@@ 6.1.2. Continuous Optimization
In the realm of continuous single-objective optimization, LLaMEA (van Stein and Bäck, 2024) utilizes LLMs to automate the evolution of algorithm design. It demonstrates the effectiveness in generating new metaphor-based optimization algorithms on BBOB benchmark (Hansen and Ros, 2010) within IOHexperimenter benchmarking tool (de Nobel et al., 2024), which supports evaluating the quality of the generated algorithms and also provides feedback to the LLM during evolution. Instead of creating new algorithms, EvolLLM (Lange et al., 2024) introduces a prompt strategy that enables LLM-based optimization to act as an Evolution Strategy (ES) and showcases robust performance on synthetic BBOB functions and neuroevolution tasks. OPRO (Yang et al., 2024) illustrates that LLMs can effectively capture optimization directions for linear regression problems by leveraging the past optimization trajectory from the meta-prompt. Additionally, LEO (Brahmachary et al., 2024) devises an explore-exploit policy using LLMs for solution generation, the method has been tested and validated in both benchmark functions as well as industrial engineering problems. Different with directly employing LLMs for generating solutions, LAEA (Hao et al., 2024) introduces LLM-based surrogate models for regression and classification tasks to assess the quality of new solutions for selection. This approach has been validated on 2D test functions using nine mainstream LLMs.
@@ 6.1.3. Bayesian Optimization
Bayesian optimization (BO) is a model-based optimization paradigm for solving expensive optimization problems and has found wide application in various real-world scenarios (Frazier and Wang, 2016). It typically employs a surrogate model to approximate the expensive function and well-designed Acquisition Functions (AFs) to select potential solutions in a sample-efficient manner. To facilitate the direct generation of solutions using LLMs, HPO-LLM* (Zhang et al., 2023) provides LLMs with an initial set of instructions that outlines the specific dataset, model, and hyperparameters to propose recommended hyperparameters for evaluation in Hyperparameter Optimization (HPO) tasks. Furthermore, LLAMBO (Liu et al., 2024a) incorporates LLM capabilities to enhance the efficiency, in which three specific enhancements throughout the BO pipeline have been systematically investigated on tasks selected from Bayesmark (Uber, 2020) and HPOBench (Eggensperger et al., 2021). Instead of utilizing LLMs for direct solution generation, BO-LIFT* (Ramos et al., 2023) utilizes predictions with uncertainties provided by a Language-Interfaced Fine-Tuning (LIFT) framework (Dinh et al., 2022) with LLMs to perform BO for catalyst optimization using natural language. EvolCAF (Yao et al., 2024a) introduces a novel paradigm to design AFs automatically for cost-aware BO, the approach showcases remarkable efficiency and generalization for synthetic functions and HPO tasks on JAHS-Bench-201 (Bansal et al., 2022). FunBO (Aglietti et al., 2024) discoveres novel and well-performing AFs for BO by extending FunSearch (Romera-Paredes et al., 2024), the discovered AFs are evaluated on various global optimization benchmarks in and out of the training distribution and HPO tasks for RBF-based SVM and AdaBoost algorithms.
@@ 6.1.4. Prompt Optimization
Prompt optimization aims to identify the most effective task prompt that maximizes the performance of the LLM on a specific task dataset. Despite of requiring specialized training for each specific task, traditional discrete or continuous approaches (Prasad et al., 2022; Li and Liang, 2021) typically necessitate access to the logits or internal states of LLMs, which may not be applicable when the LLM can only be accessed through an API. To address these issues, recent works propose to model the optimization problem in natural language with LLMs as prompts. APE (Zhou et al., 2022) utilizes the LLM as an inference model to generate instruction candidates directly based on a small set of demonstrations in the form of input-output pairs. This approach has demonstrated human-level performance on various tasks, including Instruction Induction (Honovich et al., 2022) and Big-Bench Hard (BBH) (Suzgun et al., 2022). OPRO (Yang et al., 2024) enables the LLM as an optimizer to gradually generate new prompts based on the full optimization trajectory, the optimizer prompt showcases significant improvement compared with human-designed prompts on GSM8K (Cobbe et al., 2021) and BBH. Inspired by the numerical gradient descent method, APO (Pryzant et al., 2023) conducts textual “gradient descent” by identifying the current prompts’ flaws and adjusting the prompt in the opposite semantic direction of the gradient. Similar practices are also found in the gradient-inspired LLM-based prompt optimizer named GPO (Tang et al., 2024c), as well as the collaborative optimization approach (Guo et al., 2024a) integrating a gradient-based optimizer and an LLM-based optimizer. Differently, Guo et al. (2023) introduce a discrete prompt tuning framework named EvoPrompt that prompts LLM to act like evolutionary operators in generating new candidate prompts, harnessing the benefits of evolutionary algorithms that strike a good balance between exploration and exploitation.
@@ 6.1.5. Optimization Modeling
Optimization modeling (Berry and Houston, 1995) aims to construct a mathematical model of a real-world optimization problem in a standard form, represented with an objective function and a set of constraints (AhmadiTeshnizi et al., 2023; Tang et al., 2024a). In addition to resulting significant differences in solving efficiency, creating an efficient formulation often demands expertise in optimization skills and specialized domain knowledge in various applications. Recently, LLMs have opened up promising avenues for automating optimization modeling, facilitating the formulation of problems and the implementation of solutions, thereby democratizing expert knowledge. AhmadiTeshnizi et al. (2023) propose OptiMUS, an LLM-based agent designed to formulate and solve optimization problems interpreted from natural language. The performance of the method is evaluated on a proposed novel dataset named NLP4LP (AhmadiTeshnizi et al., 2023), which includes 41 linear programming and 11 mixed integer linear programming problems. Despite its outstanding performance, the heavy reliance on sensitive data submission to proprietary LLMs can pose data privacy concerns in industrial applications.
To solve this issue, Tang et al. (2024a) propose training open-source LLMs for optimization modeling that are applicable to real-world industries. They develop and implement OR-Instruct, a semi-automated process for generating synthetic data customized to meet specific requirements. The best-performing resulting model, ORLM, demonstrates state-of-the-art performance on NL4Opt (Ramamonjison et al., 2023), MAMO (Huang et al., 2024a), and the proposed IndustryOR benchmark.
Ahmed and Choudhury (2024) develop a progressive fine-tuning framework named LM4OPT to enhance LLMs’ specificity for optimization problem formulation task on NL4Opt dataset. Wasserkrug et al. (2024) develop Decision Optimization CoPilot (DOCP), which is accessible to a broad range of decision-makers lacking problem-dependent modeling skills. This tool aids in decision-making by interacting with the user in natural language and utilizes human’s knowledge and feedback to create and solve a problem-specific optimization model. The experiments are conducted targeting three presented requirements for DOCP on facility location problems.
6.1.6. Other Applications
In addition to optimization applications that can be categorized into various domains as mentioned above, we have also discovered applications of LLMs in other optimization-related fields, which may intersect to varying degrees with the aforementioned classifications. For ease of reference, we have categorized them as ”other applications” and will introduce them separately below.
Wu et al. (2024c) propose AS-LLM, which harnesses the representation capabilities of LLMs to extract algorithm features for algorithm selection tasks. The approach integrates a feature selection module to identify critical features and leverages the similarity between algorithm and problem representations. In addition to demonstrating robust performance advantages on the ASlib (Algorithm Selection Library) benchmark (Bischl et al., 2016) in experimental studies, the authors also offer a rigorous upper bound on the model complexity of AS-LLM. Ma et al. (2024a) introduce LLaMoCo, the first instruction-tuning framework tailored to adapt general-purpose LLMs for optimizing code generation. Benefiting from a well-formatted instruction set and a novel two-phase training strategy, the fine-tuned LLMs exhibit robust and superior optimization performance, reducing the requirement for extensive domain knowledge and effort in prompt design. StrategyLLM (Gao et al., 2023) employs four LLM-based agents - the strategy generator, executor, optimizer, and evaluator - in a collaborative manner. The approach enables LLMs to reason inductively and deductively, resulting in the formulation of more generalizable and consistent few-shot prompts compared to CoT prompting methods. Experimental results demonstrate the effectiveness and reliability of StrategyLLM across four challenging tasks, encompassing math reasoning, commonsense reasoning, algorithmic reasoning, and symbolic reasoning.
@@ 6.1.6. Other Applications
EoH (Liu et al., 2024d) LLMaD FS, OS, CoT TSP, Online BPP, FSSP
@@
EvolCAF (Yao et al., 2024a) LLMaD FS, OS, CoT Synthetic Functions, JAHS-Bench-201
@@
OPRO (Yang et al., 2024) LLMaO FS GSM8K, BBH, MultiArith, AQuA
@@
StrategyLLM (Gao et al., 2023) Mixed ZS, FS, CoT Math & Commonsense & Algorithmic & Symbolic Reasoning
@@ 6.2. Machine Learning
SLINVIT (Zhang et al., 2024e) LLMaP ZS, FS ALFWorld Tasks, InterCode Tasks, BlocksWorld Tasks
@@
Neural Architecture Search EvoPrompting (Chen et al., 2023a) LLMaD FS, ZS, CoT MNIST Dataset, CLRS Algorithmic Reasoning
@@
LLM-GNN (Chen et al., 2023b) LLMaP FS, CoT Label-free Node Classification
ReStruct (Chen et al., 2024c) LLMaE FS, ZS Meta-structure Discovery
AutoAlign (Zhang et al., 2024d) LLMaE FS, ZS Entity Type Inference
@@
6.2.1. Task Planning with LLMs
6.2.2. Reinforcement Learning
Reinforcement Learning (RL) has been the de facto standard for sequential decision-making tasks, and recently, the synergy between RL and LLMs has emerged as a novel trend in the domain. This convergence mirrors the dynamics of task planning, yet places RL at the core of its methodology. Many of the LLM4AD papers on RL is for automatically designing the reward functions (Bhambri et al., 2024; Ma et al., 2024b; Narin, 2024). In addition, Shah et al. (2023) investigate the employment of LLMs for heuristic planning to steer the search process within RL frameworks. Zhang et al. (2024e) integrate LLMs into RL by introducing a Kullback-Leibler divergence regularization term that aligns LLM-driven policies with RL-derived policies. LLMs have also extended their reach to multi-agent RL scenarios, as shown by Du et al. (2024c), who illustrates their application within a Mixture-of-Experts system to direct RL models in the realm of intelligent network solutions.
6.2.3. Neural Architecture Search
6.2.4. Graph Learning
Graph learning is another application with the advancing capabilities of LLMs in symbolic reasoning and graph processing. For example, Chen et al. (2023b) apply LLMs to the task of labeling in Text-Attributed Graphs (TAGs), capitalizing on the language task proficiency of LLMs. Both Mao et al. (2024) and Chen et al. (2024c) adopt LLMs in an evolutionary framework for designing functions. The former evolves heuristic code functions to identify critical nodes in a graph while the latter identifies meta-structures within heterogeneous information networks to enhance their interpretability. Moreover, knowledge graphs have also seen substantial benefits from the application of LLMs. Zhang et al. (2024d) introduce AutoAlign, a method that employs LLMs to semantically align entities across different knowledge graphs, and Feng et al. (2023b) develop the knowledge search language to effectively conduct searches within knowledge graphs.
6.2.5. Dataset Labeling
@@ 6.3. Science Discovery
This subsection is dedicated to exploring scientific discoveries related to algorithm designs. Table. 3 summarizes the related works in this domain. For an in-depth discussion and review of LLMs in scientific discovery, readers are referred to survey papers (Birhane et al., 2023; Hadi et al., 2024; Jablonka et al., 2023).
@@ 6.3.1. General Science Discovery
In the realm of general scientific discovery, several frameworks have been developed that leverage LLMs. Notably, Du et al. (2024b) introduce LLM4ED, a framework that employs iterative strategies, including a black-box optimizer and evolutionary operators, to generate and optimize equations. This approach has shown significant advancements in stability and usability for uncovering physical laws from nonlinear dynamic systems. Similarly, Shojaee et al. (2024) present LLM-SR, which combines LLMs’ extensive scientific knowledge and code generation capabilities with evolutionary search. This framework excels in proposing and refining initial equation structures based on physical understanding, outperforming traditional symbolic regression methods in discovering physically accurate equations across multiple scientific domains. Another innovative framework, SGA, introduced by Ma et al. (2024c), is a bilevel optimization framework that merges the abstract reasoning capabilities of LLMs with the computational power of simulations. This integration facilitates hypothesis generation and discrete reasoning with simulations for experimental feedback and optimization of continuous variables, leading to improved performance in discovering constitutive laws and designing molecules.
@@ 6.3.2. Chemistry, Biology & Physics
In the field of chemistry, LLMs can be applied not only to conventional molecular generation and design (Ishida et al., 2024; Bhattacharya et al., 2024), but also to specialized areas such as drug molecule design (Ye, 2024), chemical reaction prediction and optimization (Ranković and Schwaller, 2023), and catalyst design (Wang et al., 2024a), providing customized solutions. Furthermore, LLMs have also shown promising application prospects in materials discovery, synthesis route planning, green chemistry, and other areas. These studies not only demonstrate the advantages of language models in molecular representation and optimization but also provide new tools and ideas for chemical research.
@@ 6.3.3. Mechanics
MechAgents (Ni and Buehler, 2024) introduces a class of physics-inspired generative machine learning platforms that utilize multiple LLMs to solve mechanics problems, such as elasticity, by autonomously collaborating to write, execute, and self-correct code using finite element methods. Moreover, Du et al. (2024a) adopt LLMs in automatically discovering governing equations from data, utilizing the models’ generation and reasoning capabilities to iteratively refine candidate equations. This approach, tested on various nonlinear systems including the Burgers, Chafee–Infante, and Navier–Stokes equations, demonstrates a superior ability to uncover correct physical laws and shows better generalization on unseen data compared to state-of-the-art models. Another example in fluid dynamics has been investigated by Zhu et al. (2024), which introduces FLUID-LLM, a novel framework that integrates pre-trained LLMs with spatiotemporal-aware encoding to enhance the prediction of unsteady fluid dynamics, addressing the challenge of complex geometries in computational fluid dynamics. Furthermore, Buehler (2023) study LLM-based methods for forward and inverse mechanics problems including bio-inspired hierarchical honeycomb design, carbon nanotube mechanics, and protein unfolding.

### v3

@@ 6.1.2. Continuous Optimization
In the realm of single-objective continuous optimization, LLaMEA (van Stein and Bäck, 2024) utilizes LLMs to automate the evolution of algorithm design. It demonstrates the effectiveness in generating new metaphor-based optimization algorithms on BBOB benchmark (Hansen and Ros, 2010) within IOHexperimenter benchmarking tool (de Nobel et al., 2024), which supports evaluating the quality of the generated algorithms and also provides feedback to the LLM during evolution. Instead of creating new algorithms, EvolLLM (Lange et al., 2024) introduces a prompt strategy that enables LLM-based optimization to act as an Evolution Strategy (ES) and showcases robust performance on synthetic BBOB functions and neuroevolution tasks. OPRO (Yang et al., 2024) illustrates that LLMs can effectively capture optimization directions for linear regression problems by leveraging the past optimization trajectory from the meta-prompt. Additionally, LEO (Brahmachary et al., 2024) devises an explore-exploit policy using LLMs for solution generation, the method has been tested in both benchmark functions as well as industrial engineering problems. Different with directly employing LLMs for generating solutions, LAEA (Hao et al., 2024) introduces LLM-based surrogate models for both regression and classification tasks and has been validated on 2D test functions using nine mainstream LLMs.
@@ 6.1.3. Bayesian Optimization
Bayesian optimization (BO) is a model-based optimization paradigm for solving expensive optimization problems and has found wide applications in various real-world scenarios (Frazier and Wang, 2016). It typically employs a surrogate model to approximate the expensive function and well-designed Acquisition Functions (AFs) to select potential solutions in a sample-efficient manner. To facilitate the direct generation of solutions using LLMs, HPO-LLM* (Zhang et al., 2023) provides LLMs with an initial set of instructions that outlines the specific dataset, model, and hyperparameters to propose recommended hyperparameters for evaluation in Hyperparameter Optimization (HPO) tasks. Furthermore, LLAMBO (Liu et al., 2024a) incorporates LLM capabilities to enhance BO efficiency, in which three specific enhancements throughout the BO pipeline have been systematically investigated on tasks selected from Bayesmark (Uber, 2020) and HPOBench (Eggensperger et al., 2021). Instead of utilizing LLMs for direct solution generation, BO-LIFT* (Ramos et al., 2023) utilizes predictions with uncertainties provided by a Language-Interfaced Fine-Tuning (LIFT) framework (Dinh et al., 2022) with LLMs to perform BO for catalyst optimization using natural language. EvolCAF (Yao et al., 2024a) introduces a novel paradigm integrating LLMs within EC framework to design AFs automatically for cost-aware BO, the approach showcases remarkable efficiency and generalization for synthetic functions and HPO tasks with heterogeneous costs. Similarly, FunBO (Aglietti et al., 2024) discoveres novel and well-performing AFs for BO by extending FunSearch (Romera-Paredes et al., 2024), the discovered AFs are evaluated on various global optimization benchmarks in and out of the training distribution and HPO tasks for RBF-based SVM and AdaBoost algorithms.
@@ 6.1.4. Prompt Optimization
Prompt optimization aims to identify the most effective task prompt that maximizes the performance of the LLM on a specific task dataset. Despite of requiring specialized training for each specific task, traditional discrete or continuous approaches (Prasad et al., 2022; Li and Liang, 2021) typically necessitate access to the logits or internal states of LLMs, which may not be applicable when the LLM can only be accessed through an API. To address these issues, recent works propose to model the optimization problem in natural language with LLMs as prompts. APE (Zhou et al., 2022) utilizes the LLM as an inference model to generate instruction candidates directly based on a small set of demonstrations in the form of input-output pairs. This approach has demonstrated human-level performance on various tasks, including Instruction Induction (Honovich et al., 2022) and Big-Bench Hard (BBH) (Suzgun et al., 2022). OPRO (Yang et al., 2024) enables the LLM as an optimizer to gradually generate new prompts based on the full optimization trajectory, the optimizer prompt showcases significant improvement compared with human-designed prompts on BBH and GSM8K (Cobbe et al., 2021). Inspired by the numerical gradient descent method, APO (Pryzant et al., 2023) conducts textual “gradient descent” by identifying the current prompts’ flaws and adjusting the prompt in the opposite semantic direction of the gradient. Similar practices are also found in the gradient-inspired LLM-based prompt optimizer named GPO (Tang et al., 2024c), as well as the collaborative optimization approach (Guo et al., 2024a) integrating a gradient-based optimizer and an LLM-based optimizer. Differently, Guo et al. (2023) introduce a discrete prompt tuning framework named EvoPrompt that prompts LLM to act like evolutionary operators in generating new candidate prompts, harnessing the benefits of evolutionary algorithms that strike a good balance between exploration and exploitation. StrategyLLM (Gao et al., 2023) integrates four LLM-based agents—strategy generator, executor, optimizer, and evaluator—to collaboratively induce and deduce reasoning. This method generates more generalizable and consistent few-shot prompts than CoT prompting techniques.
@@ 6.1.5. Optimization Modeling
Optimization modeling (Berry and Houston, 1995) aims to construct a mathematical model of a real-world optimization problem in a standard form, represented with an objective function and a set of constraints (AhmadiTeshnizi et al., 2023; Tang et al., 2024a). Recently, LLMs have opened up promising avenues for automating optimization modeling, facilitating the formulation of problems and the implementation of solutions, thereby democratizing expertise in optimization skills and domain knowledge in various applications. For example, OptiMUS (AhmadiTeshnizi et al., 2023) is developed as an LLM-based agent to formulate and solve optimization problems interpreted from natural language. Despite its outstanding performance on the proposed NLP4LP dataset (AhmadiTeshnizi et al., 2023), the heavy reliance on sensitive data submission to proprietary LLMs can pose data privacy concerns in industrial applications. To solve this issue, Tang et al. (2024a) propose training open-source LLMs with OR-Instruct, a semi-automated process for generating synthetic data customized to meet specific requirements. The best-performing resulting model named ORLM demonstrates state-of-the-art performance on NL4Opt (Ramamonjison et al., 2023), MAMO (Huang et al., 2024a), and the proposed IndustryOR benchmark. A related study is found in LM4OPT (Ahmed and Choudhury, 2024), which is a progressive fine-tuning framework to enhance LLMs’ specificity for optimization problem formulation task. To make it accessible to decision makers lacking problem-related modeling skills, DOCP (Wasserkrug et al., 2024) aids in decision-making by interacting with the user in natural language, and utilizes human’s knowledge and feedback to create and solve a problem-specific optimization model.
@@
EoH (Liu et al., 2024e) LLMaD FS, OS, CoT TSP, Online BPP, FSSP
@@
EvolCAF (Yao et al., 2024a) LLMaD FS, OS, CoT Synthetic Functions, HPO
@@
OPRO (Yang et al., 2024) LLMaO FS GSM8K, BBH
@@
6.1.6. Other Applications
In addition to the previously mentioned optimization applications, we have identified further uses of LLMs in related fields. For example, AS-LLM (Wu et al., 2024c) utilizes LLMs to extract and select key algorithm features for algorithm selection, enhancing the match between algorithms and problems. LLaMoCo (Ma et al., 2024a) introduces the first instruction-tuning framework for LLMs, optimizing code generation with a structured instruction set and a two-phase training approach, thereby minimizing the need for extensive domain expertise and prompt design effort.
@@ 6.2. Machine Learning
6.2.1. Task Planning with LLMs
6.2.2. Reinforcement Learning
Reinforcement Learning (RL) has been the de facto standard for sequential decision-making tasks, and recently, the synergy between RL and LLMs has emerged as a novel trend in the domain. This convergence mirrors the dynamics of task planning, yet places RL at the core of its methodology. Many of the LLM4AD papers on RL is for automatically designing the reward functions (Bhambri et al., 2024; Ma et al., 2024b; Narin, 2024). In addition, Shah et al. (2023) investigate the employment of LLMs for heuristic planning to steer the search process within RL frameworks. Zhang et al. (2024f) integrate LLMs into RL by introducing a Kullback-Leibler divergence regularization term that aligns LLM-driven policies with RL-derived policies. LLMs have also extended their reach to multi-agent RL scenarios, as shown by Du et al. (2024c), who illustrates their application within a Mixture-of-Experts system to direct RL models in the realm of intelligent network solutions.
6.2.3. Neural Architecture Search
6.2.4. Graph Learning
Graph learning is another application with the advancing capabilities of LLMs in symbolic reasoning and graph processing. For example, Chen et al. (2024b) apply LLMs to the task of labeling in Text-Attributed Graphs (TAGs), capitalizing on the language task proficiency of LLMs. Both Mao et al. (2024) and Chen et al. (2024d) adopt LLMs in an evolutionary framework for designing functions. The former evolves heuristic code functions to identify critical nodes in a graph while the latter identifies meta-structures within heterogeneous information networks to enhance their interpretability. Moreover, knowledge graphs have also seen substantial benefits from the application of LLMs. Zhang et al. (2024e) introduce AutoAlign, a method that employs LLMs to semantically align entities across different knowledge graphs, and Feng et al. (2023b) develop the knowledge search language to effectively conduct searches within knowledge graphs.
6.2.5. Dataset Labeling
@@
SLINVIT (Zhang et al., 2024f) LLMaP ZS, FS ALFWorld Tasks, InterCode Tasks, BlocksWorld Tasks
@@
Neural Architecture Search EvoPrompting (Chen et al., 2023) LLMaD FS, ZS, CoT MNIST Dataset, CLRS Algorithmic Reasoning
@@
LLM-GNN (Chen et al., 2024b) LLMaP FS, CoT Label-free Node Classification
ReStruct (Chen et al., 2024d) LLMaE FS, ZS Meta-structure Discovery
AutoAlign (Zhang et al., 2024e) LLMaE FS, ZS Entity Type Inference
@@ 6.3. Science Discovery
This subsection is dedicated to exploring LLM-based scientific discoveries related to algorithm designs. Table. 3 lists the related works in this domain.
@@ 6.3.1. General Science Discovery
In the realm of scientific discovery, LLMs are usually adopted for equation or functioin search. Notably, Du et al. (2024b) introduce LLM4ED, a framework that employs iterative strategies, including a black-box optimizer and evolutionary operators, to generate and optimize equations. This approach has shown significant advancements in stability and usability for uncovering physical laws from nonlinear dynamic systems. Similarly, Shojaee et al. (2024) present LLM-SR, which combines LLMs’ extensive scientific knowledge and code generation capabilities with evolutionary search. This framework excels in proposing and refining initial equation structures based on physical understanding, outperforming traditional symbolic regression methods in discovering physically accurate equations across multiple scientific domains. A bilevel optimization framework, named SGA, is introduce by Ma et al. (2024c). It merges the abstract reasoning capabilities of LLMs with the computational power of simulations. This integration facilitates hypothesis generation and discrete reasoning with simulations for experimental feedback and optimization of continuous variables, leading to improved performance.
@@ 6.3.2. Chemistry, Biology & Physics
In the field of chemistry, LLMs can be applied not only to conventional molecular generation and design (Ishida et al., 2024; Bhattacharya et al., 2024), but also to specialized areas such as drug molecule design (Ye, 2024), chemical reaction prediction and optimization (Ranković and Schwaller, 2023), and catalyst design (Wang et al., 2024a), providing customized solutions. Furthermore, LLMs have also shown promising application prospects in materials discovery (Jia et al., 2024; Zhang et al., 2024d), synthesis route planning (Liu et al., 2024d), green chemistry (Ruff et al., 2024), and other areas. These studies demonstrate the advantages of LLMs in molecular representation and optimization.
@@ 6.3.3. Mechanics
MechAgents (Ni and Buehler, 2024) introduces a class of physics-inspired generative machine learning platforms that utilize multiple LLMs to solve mechanics problems, such as elasticity, by autonomously collaborating to write, execute, and self-correct code using finite element methods. Moreover, Du et al. (2024a) adopt LLMs in automatically discovering governing equations from data, utilizing the models’ generation and reasoning capabilities to iteratively refine candidate equations. This approach, tested on various nonlinear systems including the Burgers, Chafee–Infante, and Navier–Stokes equations, demonstrates a superior ability to uncover correct physical laws and shows better generalization on unseen data compared to other models. Another example in fluid dynamics has been investigated by Zhu et al. (2024), which introduces FLUID-LLM, a novel framework that integrates pre-trained LLMs to enhance the prediction of unsteady fluid dynamics. AutoTurb (Zhang et al., 2024g), on the other hand, adopts LLMs in an evolutionary framework to automatically design and search for turbulence model in computational fluid dynamics. Furthermore, Buehler (2023) study LLM-based methods for forward and inverse mechanics problems including bio-inspired hierarchical honeycomb design, carbon nanotube mechanics, and protein unfolding.
@@
AutoTurb (Zhang et al., 2024g) LLMaD Computational Fliud Dynamics